OpenStudy (anonymous):

what is the equation of a line that is perpendicular to the line whose equation is y=x+2?

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

Find a point first on the line \[y=x+2\] The most easiest point is either the x or y intercept.

OpenStudy (anonymous):

Let's just pick the x intercept. That means you just let x=0 and then you can find the y-coordinate for the x-intercept. So the x-intercept would be the point (0, y-coordinate)

OpenStudy (anonymous):

y = -x + k. k can be any value. If we have to find the line perpendicular to y=x+2 and passing through some point then we can easily solve k.

OpenStudy (anonymous):

so whats our anwer then

OpenStudy (anonymous):

y = -x + k. for each k one line will exist. So there are infinite number of parallel lines exist those are perpendicular to y = x+2

OpenStudy (anonymous):

@sriramreddy please show her how to do it by herself and restrain yourself from giving the answer away until she knows how to do it.